OpenStudy (eric_d):

Express as the sum or difference of two trigonometric ratios sin(theta + 75)sin(theta + 15) @ganeshie8

OpenStudy (eric_d):

@ikram002p

ganeshie8 (ganeshie8):

Lookup 2sinAsinB formula

OpenStudy (ikram002p):

lol i dint get it yet

OpenStudy (eric_d):

@ganeshie8

ganeshie8 (ganeshie8):

\[\large 2 \sin A \sin B = \cos(A-B) - \cos (A+B)\]

ganeshie8 (ganeshie8):

\[\large \sin(\theta + 75)\sin(\theta + 15) = \dfrac{1}{2}*2 \sin(\theta + 75)\sin(\theta + 15) \\ \ \\\large = ? \]

ganeshie8 (ganeshie8):

Notice that \(A = \theta + 75\) \(B = \theta + 15\)

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{2}*2 \sin(\theta + 75)\sin(\theta + 15) = ? \]

ganeshie8 (ganeshie8):

\(A - B = 60\) \(A + B = 2\theta + 90\)

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{2}*2 \sin(\theta + 75)\sin(\theta + 15) = \dfrac{1}{2}\left[\cos (60) - \cos (2\theta + 90)\right] \]

ganeshie8 (ganeshie8):

see if that looks okay so far

OpenStudy (eric_d):